OpenStudy (anonymous):

need the solutions x^2-3x-1=0

OpenStudy (anonymous):

\[\frac{-(-3)\pm \sqrt{3^3-4*(-1)}}{2}\]

OpenStudy (anonymous):

\[\frac{3\pm \sqrt{9+4}}{2}=\frac{3\pm \sqrt{13}}{2}\]